# Belinda Z. Li    Zifan Carl Guo    Jacob Andreas

### v2

@@ Abstract
Transformer language models (LMs) exhibit behaviors—from storytelling to code generation—that appear to require tracking the unobserved state of an evolving world. How do they do so? We study state tracking in LMs trained or fine-tuned to compose permutations (i.e., to compute the order of a set of objects after a sequence of swaps). Despite the simple algebraic structure of this problem, many other tasks (e.g., simulation of finite automata and evaluation of boolean expressions) can be reduced to permutation composition, making it a natural model for state tracking in general. We show that LMs consistently learn one of two state tracking mechanisms for this task. The first closely resembles the “associative scan” construction used in recent theoretical work by Liu et al. (2023) and Merrill et al. (2024). The second uses an easy-to-compute feature (permutation parity) to partially prune the space of outputs, then refines this with an associative scan. The two mechanisms exhibit markedly different robustness properties, and we show how to steer LMs toward one or the other with intermediate training tasks that encourage or suppress the heuristics. Our results demonstrate that transformer LMs, whether pretrained or fine-tuned, can learn to implement efficient and interpretable state tracking mechanisms, and the emergence of these mechanisms can be predicted and controlled. [^1] Code and data are available at https://github.com/belindal/state-tracking
@@
Figure 1: We use permutation word problems as a simple model of state tracking. Here actions permutations, and states are the products of those permutations; the current state can be tracked by taking the cumulative product from left to right (§2). We identify several possible algorithms that Transformers may use to solve permutation word problems, which we call sequential, parallel, associative, and parity-associative (§3). Above, we depict the “signatures” of each algorithm under two types of interpretability analysis: prefix patching, where all the activations are corrupted except the prefix up to a token at a particular layer, and probing, where we train a linear probe to map from last-token representations across the layers to either the final state or the final state parity (§2.3).
@@ 1 Introduction
Language models (LMs) are trained to model the surface form of text. A growing body of work suggests that these models nevertheless learn to represent the latent state of the world—e.g. situations described by language and results of program execution—to support prediction. However, the mechanisms that LMs use to construct these representations are not understood. Do LMs simulate state evolution step-by-step across successive hidden layers or token representations (Yang et al., 2024)? Are states approximated via a complex collection of heuristics (jylin04 et al., 2024)? Is state tracking an illusion (Bender & Koller, 2020)?
@@
This paper studies the implementation and emergence of state tracking mechanisms in language models using permutation composition as a model system: given a fixed set of objects, we train or fine-tune LMs to predict the final position of each object after a sequence of rearrangements. Previous work has used versions of this task to evaluate LMs’ empirical state tracking abilities (Li et al., 2021; Kim & Schuster, 2023; Li et al., 2023). Additionally, as shown by Barrington (1986), many complex, natural state tracking tasks—including simulation of finite automata and evaluation of Boolean expressions—can be reduced to permutation tracking with five or more objects. This makes it a natural model for studying state tracking in general.
@@
Our analysis proceeds in several steps. Section 2 provides technical preliminaries: §2.1 and §2.2 introduce state tracking problems and the permutation composition task we use to model them (Figure 1A), and §2.3 reviews the set of interpretability tools we use to analyze LM computations. Next, Section 3 lays out a family of algorithms that past work has suggested LMs might, in principle, use to solve the state tracking task (Figure 1D), and describes the signatures—expected readouts from different interpretability methods—that we would expect to find if a given algorithm is implemented (Figure 1B-C).
@@
Finally, Sections 4 and 5 present experimental findings. Across a range of sizes, architectures, and pretraining schemes, we find that LMs consistently learn one of two state tracking mechanisms. The first mechanism, which we call the “associative algorithm” (AA), resembles the associative scan construction used by Liu et al. (2023) and Merrill & Sabharwal (2024) to establish theoretical lower bounds on the expressive capacity of Transformers. The second mechanism, which we call the “parity-associative algorithm” (PAA), first rules out a subset of final states using an easy-to-compute permutation parity heuristic, then uses an associative scan to obtain a final-state prediction. Notably, we fail to find evidence for either step-by-step simulation or for fully parallel composition mechanisms, even when these are theoretically implementable by LMs. We support our findings with evidence from representation interventions (Meng et al., 2022; Zhang & Nanda, 2024; §4.2), probes (Shi et al., 2016; §4.3), patterns in prediction errors (Zhong et al., 2024; (§4.4), attention maps (Clark et al., 2019; §4.5), and training dynamics (McCoy et al., 2019; Olsson et al., 2022; Hu et al., 2023; §5.1).
@@
As pretrained LMs sometimes re-use circuits when fine-tuned on related tasks (Prakash et al., 2024; Merullo et al., 2024), our results suggest a possible mechanism by which real-world LMs might perform state tracking when modeling language, code, and games. Looking beyond state tracking, these findings underscore both the complexity and variability of LM solutions to complex tasks, which may involve both heuristic features and structured solutions.
@@ 2.1 State Tracking
Inferring common ground in discourse (Li et al., 2021), navigating the environment (Vafa et al., 2024), reasoning about code (Merrill et al., 2024), and playing games (Li et al., 2023; Karvonen, 2024) all require being able to track the evolving state of a real or abstract world. There has been significant interest in understanding whether (and how) LMs can perform these tasks. In theoretical work, researchers have observed that many natural state-tracking problems (including the ones listed above) are associated with the complexity class $\mathsf{NC}^{1}$, and have shown that inputs of size $n$ can be modeled by Transformers with $O(\log n)$ depth (Liu et al., 2023; Merrill & Sabharwal, 2024). Empirical work, meanwhile, has found that large LMs learn to solve state tracking problems (Kim & Schuster, 2023) and encode state information in their representations (Li et al., 2021; Li et al., 2023). But a mechanistic understanding of how trained LMs infer these states has remained elusive.
@@ 2.2 Permutation Group Word Problems
The composition of two permutations, standardly denoted $a_{1}\circ a_{0}$, is the result of applying $a_{1}$ after $a_{0}$. Inputs to sequence models in machine learning are typically written with earlier inputs before later inputs (i.e. left-to-right), so for consistency with this convention we will write $a_{0}a_{1}$ to denote the application of $a_{0}$ then $a_{1}$. Figure 1A shows the result of composing 42315 and 12534 in sequence.
Finally, the word problem on $\boldsymbol{S}_{n}$ is the problem of computing the product of a sequence of permutations. This product itself corresponds to a single permutation (32514 in Figure 1). But, following the intuition given at the beginning of the section, it may equivalently be interpreted as the final ordering of the objects being rearranged (DBAEC in Figure 1A). In accordance with this intuitive explanation (and by analogy to other state tracking problems), we will use $a_{t}$ to denote a single permutation (“action”) in a sequence, and $s_{t}=a_{0}\cdots a_{t}$ to denote the result of a sequence of permutations (a “state”).
@@
All experiments in this paper train transformer language models to solve the word problem: they take as input a sequence of actions $[a_{0},\ldots,a_{t}]$, and output a sequence of state predictions $[s_{0},\ldots,s_{t}]$.
@@ 2.3 Interpretability Methods
In probing experiments (Shi et al., 2016), we fix the target LM, then train a smaller “probe” model (e.g. a linear classifier) to map LM hidden representations $h$ to quantities $z$ hypothesized to be encoded by the LM (Figure 1C). Our experiments specifically evaluate whether (1) the state $s_{t}$, and (2) the final state parity is linearly encoded in intermediate-layer representations. For each layer $l$, we train (1) a state probe to predict $p(s_{t}\mid h_{t,l})$ ; and (2) a parity probe to predict $p(\epsilon(s_{t})\mid h_{t,l})$. Given a trained LM, we collect representations on one set of input sequences to train the probe, then evaluate probe accuracy on a held-out set.
@@
We then re-run the LM on the corrupted input $x^{\prime}$, but substitute a hidden representation from the clean input $x$, and measure how much prediction shifts toward the clean output $\hat{y}$ using the normalized logit difference (Wang et al., 2022):
@@
In this paper, we evaluate which representations are involved in prediction by presenting models with a clean sequence $[a_{0},a_{1},\ldots,a_{t}]$ associated with a final state $s_{t}$. We then produce a corrupted sequence $[a_{0}^{\prime},a_{1},\ldots,a_{t}]$, associated with a final state $s_{t}^{\prime}$, and identify the hidden states that, when patched in, cause the model to output $s_{t}$ rather than $s_{t}^{\prime}$ with high probability. Our main experiments specifically perform prefix patching, where all hidden representations up to index $t$ ($h_{1:t,l}\leftarrow h_{1:t,l}^{\text{clean}}$) are patched at a particular layer $l$ (Figure 1B). Prefix patching allows us to localize how information gets progressively transferred to the final token as we move down the network. An NLD value close to 1 means that some part of the prefix representation was used for prediction, while a value close to 0 means that no part was.
@@ 3 What Algorithms can Transformers Implement in Theory?
To use the methods described in Section 2.3 to interpret model behavior, we must first establish a phenomenology for LM state tracking—identifying candidate state tracking algorithms that might be implemented by the model, along with the empirical probing and activation patching results we would expect to find if these algorithms are implemented. Below, we describe a set of state tracking mechanisms suggested by the existing literature.
@@ 3.1 Sequential Algorithm
The sequential algorithm composes permutations one at a time from left to right (analogous to a mechanism some LMs use to solve multi-hop reasoning problems; Yang et al., 2024). Signatures of this algorithm would provide evidence that LMs implement step-by-step “simulation” in their hidden states to solve state tracking tasks. In this algorithm, each hidden state $h_{t,l}$ stores the associated action $a_{t}$ until $s_{t}$ can be computed, maintaining $h_{t,t}=s_{t}$.
@@
The first row of Figure 1D shows the set of data dependencies between states when using this algorithm to compute the final state in a sequence. It can be seen (both from this figure and the algorithm above) that this computation depends only on hidden states with $l\leq t$.
@@
Because of this dependency, any patching experiment that replaces only hidden states with $l>t$ will have no effect on the final model predictions, leading to the upper triangular patching signature shown in the first row of Figure 1B.
@@
Because $s_{t}$ can only be predicted at layer $l=t$, we expect a state probe to show a linear dependence on depth: for sequences of maximum length $T$, a probe at layer $l$ will correctly label an $l/T$ fraction of states. If these state representations linearly encode parity, then the accuracy of the parity probe will also increase linearly, otherwise it will remain constant. [^2] See Section C.1 for a representative model in which parity is not linearly decodable.
@@ 3.2 Parallel Algorithm
As noted in Section 2.2, the word problem on $S_{5}$ belongs to $\mathsf{NC}^{1}$ (and thus requires a circuit depth that scales logarithmically with sequence length). The word problem on $S_{3}$, however, belongs to $\mathsf{TC}^{0}$, the class of decision problems with constant-depth threshold circuits. See discussion in Appendix A and Merrill & Sabharwal (2023). A constant-depth circuit will give rise to a set of hidden-state dependencies like the second row of Figure 1D.
@@ 3.3 Associative Algorithm
In the associative algorithm (AA), Transformers compose permutations hierarchically: in each layer, adjacent sequences of permutations are grouped together and their product is computed. This is analogous to recursive scan in Liu et al. (2023) and flattened expression evaluation in Merrill et al. (2024). This algorithm takes advantage of the associative nature of the product of permutations, whereby $a_{0}a_{1}a_{2}a_{3}=(a_{0}a_{1})(a_{2}a_{3})$. It ensures that $h_{t,l}=a_{t-2^{l}+1}\cdots a_{t}$, and thus that $h_{t,\log(t+1)}=a_{0}\cdots a_{t}$. Signatures of this algorithm would provide evidence that LMs perform state tracking not by encoding states, bur rather maps from initial to final states, for prefixes of increasing length.
@@
As can be seen from the dataflow visualization in the third row of Figure 1D, the model’s prediction for $s_{t}$ depends on the hidden representation $h_{t/2}$ in the layer before the final state is computed, the representation at $h_{t/4}$ in the layer before that, etc.
@@
Consequently, for AA, the length of the prefix that must be modified to alter model behavior increases exponentially in depth, resulting in the signature seen in the third row of Figure 1B.
@@ 3.4 Parity-Associative Algorithm
In this algorithm (PAA), LMs compute the final state in two stages: first computing the parity of the state (which can be performed in a constant number of layers using a subroutine from the Parallel algorithm); then separately computing the remaining information needed to identify the final state (the “parity complement”) using a procedure analogous to AA. (Unlike the preceding algorithms, we are not aware of any previous proposals for solving permutation composition problems in this way—but as we will see, it is useful for understanding interactions between “heuristic” and “algorithmic” solutions in real LMs.)
@@
We expect state probes to improve exponentially with depth, while parity probes converges to 100% at a constant depth.

### v3

@@ Abstract
Transformer language models (LMs) exhibit behaviors—from storytelling to code generation—that seem to require tracking the unobserved state of an evolving world. How do they do this? We study state tracking in LMs trained or fine-tuned to compose permutations (i.e., to compute the order of a set of objects after a sequence of swaps). Despite the simple algebraic structure of this problem, many other tasks (e.g., simulation of finite automata and evaluation of boolean expressions) can be reduced to permutation composition, making it a natural model for state tracking in general. We show that LMs consistently learn one of two state tracking mechanisms for this task. The first closely resembles the “associative scan” construction used in recent theoretical work by Liu et al. (2023) and Merrill et al. (2024). The second uses an easy-to-compute feature (permutation parity) to partially prune the space of outputs, and then refines this with an associative scan. LMs that learn the former algorithm tend to generalize better and converge faster, and we show how to steer LMs toward one or the other with intermediate training tasks that encourage or suppress the heuristics. Our results demonstrate that transformer LMs, whether pre-trained or fine-tuned, can learn to implement efficient and interpretable state-tracking mechanisms, and the emergence of these mechanisms can be predicted and controlled. Code and data are available at https://github.com/belindal/state-tracking.
@@
Figure 1: We use permutation word problems as a simple model of state tracking. Actions are permutations, and states are the products of those permutations; the current state can be tracked by taking the cumulative product from left to right (§2). We identify several possible algorithms that Transformers may use to solve permutation word problems: sequential, parallel, associative, and parity-associative (§3). Above, we depict the “signatures” of each algorithm under two types of interpretability analysis: prefix patching, where we create pairs of prompts differing only on the first token, then substitute all activation except the prefix up to a token at a particular layer, and probing, where we train a linear probe to map from last-token representations across the layers to either the final state or the final state parity (§2.3). Note: the dotted lines indicate two different probing signatures consistent with this algorithm (see Section C.3 for more details).
@@ 1 Introduction
Language models (LMs) are trained to model the surface form of text. A growing body of work suggests that model internals contain a latent, decodable state of the world—e.g., situations described by language and results of program execution—to support prediction (Li et al., 2021; Nanda et al., 2023; Li et al., 2023). However, the mechanisms that LMs use to construct these representations are not understood. Do LMs simulate state evolution step by step across successive hidden layers or token representations (Yang et al., 2024)? Are states approximated through a complex collection of heuristics (jylin04 et al., 2024)? Is state tracking an illusion (Bender & Koller, 2020)?
@@
This paper studies the implementation and emergence of state tracking mechanisms in language models using permutation composition as a model system: given a fixed set of objects, we train or fine-tune LMs to predict the final position of each object after a sequence of rearrangements. Previous work has used versions of this task to evaluate LMs’ empirical state tracking abilities (Li et al., 2021; Kim & Schuster, 2023; Li et al., 2023). Additionally, as shown by Barrington (1989), many complex, natural, state-tracking tasks—including simulation of finite automata and evaluation of Boolean expressions—can be reduced to permutation tracking with five or more objects. This makes it a natural model for studying state tracking in general.
@@
Our analysis proceeds in several steps. §2 provides technical preliminaries: §2.1 and §2.2 introduce state tracking problems and the permutation composition task we use to model them (Figure 1A), and §2.3 reviews the set of interpretability tools we use to analyze LM computations. Next, §3 lays out a family of algorithms that past work has suggested LMs might, in principle, use to solve the state tracking task (Figure 1D), and describes the signatures—expected readouts from different interpretability methods—that we would expect to find if a given algorithm is implemented (Figure 1B-C).
@@
Finally, §4 and §5 present experimental findings. Across a range of sizes, architectures, and pretraining schemes, we find that LMs consistently learn one of two state tracking mechanisms. The first mechanism, which we call the “associative algorithm” (AA), resembles the associative scan construction used by Liu et al. (2023) and Merrill & Sabharwal (2024) to establish theoretical lower bounds on the expressive capacity of Transformers. The second mechanism, which we call the “parity-associative algorithm” (PAA), first rules out a subset of final states using an easy-to-compute permutation parity heuristic, then uses an associative scan to obtain a final prediction. Notably, we fail to find evidence for either step-by-step simulation or fully parallel composition, despite their being theoretically implementable by LMs. We support our findings with evidence from representation interventions (Meng et al., 2022; Zhang & Nanda, 2024; §4.2), probes (Shi et al., 2016; §4.3), patterns in prediction errors (Zhong et al., 2024; §4.4), attention maps (Clark et al., 2019; §4.5), and training dynamics (McCoy et al., 2020; Olsson et al., 2022; Hu et al., 2023; §5.1).
@@
As pretrained LMs sometimes re-use circuits when fine-tuned on related tasks (Prakash et al., 2024; Merullo et al., 2024), our results suggest a possible mechanism by which real-world LMs might perform state tracking when modeling language, code, and games. We show preliminary evidence of these algorithms on a version of our permutation composition tasks expressed in natural language (Appendix E). Looking beyond state tracking, these findings underscore both the complexity and variability of LM solutions to complex tasks, which may involve both heuristic features and structured solutions.
@@ 2.1 State Tracking
Inferring common ground in discourse (Li et al., 2021), navigating the environment (Vafa et al., 2024), reasoning about code (Merrill et al., 2024), and playing games (Li et al., 2023; Karvonen, 2024) all require being able to track the evolving state of a real or abstract world. There has been significant interest in understanding whether (and how) LMs can perform these tasks. In theoretical work, researchers have observed that many natural state-tracking problems (including the ones listed above) are associated with the complexity class $\mathsf{NC}^{1}$, but Transformers cannot track the state of arbitrarily long inputs (Merrill et al., 2022; Huet et al., 2025; Bhattamishra et al., 2020; Delétang et al., 2023; Strobl et al., 2024). However, prior work has shown that Transformers with $O(\log n)$ depth can model inputs of up to length $n$ (Liu et al., 2023; Merrill & Sabharwal, 2024). Empirical work, meanwhile, has found that large LMs learn to solve state tracking problems (Kim & Schuster, 2023) and encode state information in their representations (Li et al., 2021; Li et al., 2023). But a mechanistic understanding of how trained LMs infer these states has remained elusive.
@@ 2.2 Permutation Group Word Problems
The composition of two permutations, standardly denoted $a_{1}\circ a_{0}$, is the result of applying $a_{1}$ after $a_{0}$. Inputs to sequence models in machine learning are typically written with earlier inputs before later inputs (i.e., left-to-right), so for consistency with this convention, we will write $a_{0}a_{1}$ to denote the application of $a_{0}$ then $a_{1}$. Figure 1A shows the result of composing 42315 and 12534 in sequence.
Finally, the word problem on $\boldsymbol{S}_{n}$ is the problem of computing the product of a sequence of permutations. This product itself corresponds to a single permutation (32514 in Figure 1). But, following the intuition given at the beginning of the section, it may equivalently be interpreted as the final ordering of the objects being rearranged (DBAEC in Figure 1A). Following this intuitive explanation (and by analogy to other state tracking problems), we will use $a_{t}$ to denote a single permutation (“action”) in a sequence, and $s_{t}=a_{0}\cdots a_{t}$ to denote the result of a sequence of permutations (a “state”).
@@
All experiments in this paper train transformer language models to solve the word problem: they take as input a sequence of actions $a_{0},\ldots,a_{t}]$, and output a sequence of state predictions $s_{0},\ldots,s_{t}]$. We also validate our findings on a natural language version of this task in Appendix E, where permutations are expressed as instructions like swap positions 2 and 3.
@@ 2.3 Interpretability Methods
In probing experiments (Shi et al., 2016), we fix the target LM, then train a smaller “probe” model (e.g. a linear classifier) to map LM hidden representations $h$ to quantities $z$ hypothesized to be encoded by the LM (Figure 1C). Our experiments specifically evaluate whether (1) the state $s_{t}$, and (2) the final state parity is linearly encoded in intermediate-layer representations. For each layer $l$, we train (1) a state probe to predict $p(s_{t}\mid h_{t,l})$and (2) a parity probe to predict $p(\epsilon(s_{t})\mid h_{t,l})$. Given a trained LM, we collect representations on one set of input sequences to train the probe, then evaluate probe accuracy on a held-out set.
@@
We then re-run the LM on the corrupted input $x^{\prime}$, but substitute a hidden representation from the clean input $x$, and measure how much prediction shifts toward the clean output $\hat{y}$ using the normalized logit difference (Wang et al., 2023):
@@
In this paper, we evaluate which representations are involved in prediction by presenting models with a clean sequence $a_{0},a_{1},\ldots,a_{t}]$ associated with a final state $s_{t}$. We then produce a corrupted sequence differing only in the first token, $a_{0}^{\prime},a_{1},\ldots,a_{t}]$, associated with a final state $s_{t}^{\prime}$. We then identify the hidden states that, when patched in, cause the model to output $s_{t}$ rather than $s_{t}^{\prime}$ with high probability. Our main experiments specifically perform prefix patching, where all hidden representations up to index $t$ ($h_{1:t,l}\leftarrow h_{1:t,l}^{\text{clean}}$) are patched at a particular layer $l$ (Figure 1B). Prefix patching allows us to localize how information gets progressively transferred to the final token as we move deeper into the network. A value close to 1 means that some part of the prefix representation was used for prediction; a value close to 0 means that no part was.
@@ 3 What Algorithms Can Transformers Implement in Theory?
To use the methods described in §2.3 to interpret model behavior, we must first establish a phenomenology for LM state tracking—identifying candidate state tracking algorithms that might be implemented by the model, along with the empirical probing and activation patching results we would expect to find if these algorithms are implemented. Below, we describe a set of state tracking mechanisms suggested by the existing literature.
@@ 3.1 Sequential Algorithm
The sequential algorithm composes permutations one at a time from left to right (analogous to a mechanism some LMs use to solve multi-hop reasoning problems; Yang et al., 2024). Signatures of this algorithm would provide evidence that LMs implement step-by-step “simulation” in their hidden states to solve state tracking tasks. In this algorithm, each hidden state $h_{t,l}$ stores the associated action $a_{t}$ until $s_{t}$ can be computed, maintaining $h_{t,t}=s_{t}$. As shown in the first row of Figure 1D, this computation depends only on hidden states with $l\leq t$.
@@
Because of this dependency, any patching experiment that replaces only hidden states with $l>t$ will not affect the final model predictions, leading to the upper triangular patching signature shown in the first row of Figure 1B.
@@
Because $s_{t}$ can only be predicted at layer $l=t$, we expect a state probe to show a linear dependence on depth: for sequences of maximum length $T$, a probe at layer $l$ will correctly label an $l/T$ fraction of states. If these state representations linearly encode parity, then the accuracy of the parity probe will also increase linearly; otherwise, it will remain constant. [^1] See Section C.3 for a representative model in which parity is not linearly decodable.
@@ 3.2 Parallel Algorithm
As noted in §2.2, the word problem on $S_{5}$ belongs to $\mathsf{NC}^{1}$ (and thus requires a circuit depth that scales logarithmically with sequence length). The word problem on $S_{3}$, however, belongs to $\mathsf{TC}^{0}$, the class of decision problems with constant-depth threshold circuits. See discussion in Appendix A and Merrill & Sabharwal (2023). A constant-depth circuit will give rise to a set of hidden-state dependencies like the second row of Figure 1D.
@@ 3.3 Associative Algorithm
In the associative algorithm (AA), Transformers compose permutations hierarchically: in each layer, adjacent sequences of permutations are grouped together and their product is computed. This is analogous to recursive scan in Liu et al. (2023) and flattened expression evaluation in Merrill et al. (2024). This algorithm takes advantage of the associative nature of the product of permutations, whereby $a_{0}a_{1}a_{2}a_{3}=(a_{0}a_{1})(a_{2}a_{3})$. It ensures that $h_{t,l}=a_{t-2^{l}+1}\cdots a_{t}$, and thus that $h_{t,\log(t+1)}=a_{0}\cdots a_{t}$. Signatures of this algorithm would provide evidence that LMs perform state tracking not by encoding states, but rather by mapping between states, for prefixes of increasing length.
@@
As seen in the third row of Figure 1D, the model’s prediction for $s_{t}$ depends on the hidden representation $h_{t/2}$ in the layer before the final state is computed, the representation at $h_{t/4}$ in the layer before that, etc.
@@
Consequently, for AA, the length of the prefix that must be modified to alter model behavior increases exponentially in depth, resulting in the signature in the third row of Figure 1B.
@@ 3.4 Parity-Associative Algorithm
In this algorithm (PAA), LMs compute the final state in two stages: first computing the parity of the state (which can be performed in a constant number of layers using a subroutine from the Parallel algorithm); then separately computing the remaining information needed to identify the final state (the “parity complement”) using a procedure analogous to AA. (Unlike the preceding algorithms, we are not aware of any previous proposals for solving permutation composition problems in this way; but as we will see, it is useful for understanding interactions between “heuristic” and “algorithmic” solutions in real LMs.)
@@
We expect state probes to improve exponentially with depth, while parity probes converge to 100% at a constant depth.